OpenStudy (anonymous):

EASY HELP: PLEASE: A bag contains two red marbles, three green ones, one lavender one, two yellows, and three orange marbles. How many sets of four marbles include all the red ones?

OpenStudy (anonymous):

i thought it was C(3,3)=1 C(9,1)=9 so i thought the answer was 9 but for some reason its wrong

OpenStudy (anonymous):

oh hey @ganeshie8 :)

ganeshie8 (ganeshie8):

okay, so u wanto to choose 4 marbles

ganeshie8 (ganeshie8):

two marbles are are already choosen by god for u : 2 REDS are a must

OpenStudy (anonymous):

sorry 2,2

ganeshie8 (ganeshie8):

that means you need to choose 2 more marbles

OpenStudy (anonymous):

hmm so 2 more?

ganeshie8 (ganeshie8):

Count total number of marbles first

OpenStudy (anonymous):

11 total

ganeshie8 (ganeshie8):

2R 3G 1L 2Y 3O

ganeshie8 (ganeshie8):

yes, 11 total. and since 2 are in ur set already : you need to pick the remaining two from 11-2 = 9 marbles

ganeshie8 (ganeshie8):

how many ways u can choose 2 things from 9 ?

OpenStudy (anonymous):

7

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

4.5?

ganeshie8 (ganeshie8):

3G 1L 2Y 3O : 2G, 1G, 1L 1G, 1Y 1G, 1O 1L, 1Y 1L, 1O 2Y, 1Y, 1O 2O

ganeshie8 (ganeshie8):

9 should be the correct answer right ?

ganeshie8 (ganeshie8):

did the grader marked it wrong ?

OpenStudy (anonymous):

yes but for some reason it is wrong

OpenStudy (anonymous):

might be a glitch in the system?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

let me think a bit more..

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

9 should be correct

ganeshie8 (ganeshie8):

@iambatman wat do u think

OpenStudy (anonymous):

yeah exactly hmm

ganeshie8 (ganeshie8):

maybe try entering 36 once

OpenStudy (anonymous):

yeah it was 36 hmm but why not 9?

ganeshie8 (ganeshie8):

cool :) they want u treat each marble as different : think of the 3 green marbles as : {G1, G2, G3}

ganeshie8 (ganeshie8):

dont think of them as : {G, G, G}

OpenStudy (anonymous):

oh so you multiply them all?

ganeshie8 (ganeshie8):

3G 1L 2Y 3O : then total number of sets having 2R and 2G marbles would be : \(\large^3C_2 \)

ganeshie8 (ganeshie8):

cuz, u can choose 2 Green marbles from 3 Green marbles in : \(\large^3C_2\) ways

OpenStudy (anonymous):

and C is the total added together correct

ganeshie8 (ganeshie8):

yes you need to add up all cases

ganeshie8 (ganeshie8):

check the wolfram link :)

OpenStudy (anonymous):

ok awesome thank you! :) @ganeshie8

ganeshie8 (ganeshie8):

2 Green : \(\large ^3C_2\) 1 Green 1 Lavender : \(\large ^3C_1 \times ^1C_1\) 1 Green 1 Lavender : \(\large ^3C_1 \times ^2C_1\) 1 Green 1 Orange : \(\large ^3C_1 \times ^3C_1\) .......

ganeshie8 (ganeshie8):

add up all the cases... wolfram link shows the exact same

OpenStudy (anonymous):

ok awesome! @ganeshie8

ganeshie8 (ganeshie8):

np :) @mathfriend any other shortcut u have other than the dumb method i suggested ?

OpenStudy (anonymous):

i think there might be a shortcut @ganeshie8 i can't think of it at the moment lol

OpenStudy (anonymous):

Basically, there are 2 red marbles, and 9 other marbles. So you're looking at how many combinations can these 9 marbles make by taking 2 at a time. 9*8/2 = 36...Does this make sense?

OpenStudy (anonymous):

The wording to this question is terrible, that's why it's tricky :p

ganeshie8 (ganeshie8):

that makes perfect sense now after looking at it again xD

OpenStudy (anonymous):

it does!

OpenStudy (anonymous):

i hate the wording at times haha

OpenStudy (anonymous):

Haha, sorry for the late reply, I kept thinking I was wrong...so I had to keep rereading the damn question.

OpenStudy (anonymous):

no worries its ok its cool to learn both ways !!

OpenStudy (anonymous):

Alright, time for me to go fight crime, g'night :p

ganeshie8 (ganeshie8):

thanks @iambatman , u deserve a good night's sleep ! embrace urself tightly and sleep :) good night !!

OpenStudy (anonymous):

haha awesome good night!!!

OpenStudy (anonymous):

Thank you @ganeshie8 good night, and take care! @hogwardsacademy Good night, and good luck!

OpenStudy (anonymous):

Thank you! @iambatman

OpenStudy (anonymous):

Hmmmm..... I would have said a required set would look something like {R,R,G,L} or {O,Y,R,R} or {R,O,O,R} etc. and its how many of these you can have. Lets look at the sets whereby the two reds come at the start and lavendar - if present at all - comes at the end. Then you have the following distinct 9 sets: { R, R, G, G } { R, R, Y, Y } { R, R, O, O } { R, R, G, Y } { R, R, G, O } { R, R, G, L } { R, R, Y, O } { R, R, Y, L } { R, R, O, L } So if the order was not important, the answer would be 9. If the order was relevant then then each of the first 3 sets can be arranged 6 unique ways. Each of the remaining 6 can be re-organised into 12 distinct sets - giving 3 x 6 + 6 x 12 = 90 So I would have said 90. But I guess it comes down to the interpretation of the question.

ganeshie8 (ganeshie8):

I see, yeah the problem must be formulated a bit more clearly...